Mr. Hernandez has 1/3 of his saving in mutual funds, 1/4 of his savings in bonds, and the rest of his savings in his bank account. What is the total amount of savings if he has $30,000 in the bank account?
Answer by richwmiller(17219)

1-1/3-1/4=b
b=5/12
5/12*s=30000
s = 72000